OpenStudy (anonymous):

SIMPLIFY! ~ 6 with the exponent of 3 over 6

hero (hero):

6^3/6 = 216/6 = 36

OpenStudy (anonymous):

Or you can go at it this way \[\frac{6^3}{6}=\frac{6^3}{6^1}\]\[=6^{3-1}\]\[=6^2\]\[=36\]

OpenStudy (anonymous):

thanks <333333333333333